A discrete random variable has the following distribution
X 1 2 3 4
f(x) k 2k 2/3 k 1/3 k

i).Determine the value of k (2mks)
ii).Find the expectation of X (2mks)
iii).Find the variance of X
1
Expert's answer
2020-03-11T14:17:15-0400

i) "\Sigma p_i=1"

"k+2k+\frac{2}{3}k+\frac{1}{3}k=1"

"4k=1, k=\frac{1}{4}"


ii) "E(X)=\Sigma x_ip_i=1*\frac{1}{4}+2*\frac{2}{4}+3*\frac{2}{12}+4*\frac{1}{12}=\frac{25}{12}"


iii)

"Var(X)=\Sigma x_i^2 p_i-(E(X))^2=1^2*\frac{1}{4}+2^2*\frac{2}{4}+3^2*\frac{2}{12}+\\+4^2*\frac{1}{12}-(\frac{25}{12})^2=\frac{107}{144}"